OpenStudy (anonymous):

if y =(1-u)/(1+u) and u = cos 2x, what is y' ? a.2tanx sec ^2 x b.2secx tan^2 x c.-2tan x sec^2 x d.-2secx tan^2 x

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{dy}{du}\frac{du}{dx}\] is what you need

OpenStudy (anonymous):

first one is the right one